# MINGLE: VLMs for Semantically Complex Region Detection in Urban Scenes

### v2

@@ Abstract
Understanding group-level social interactions in public spaces is crucial for urban planning, informing the design of socially vibrant and inclusive environments. Detecting such interactions from images involves interpreting subtle visual cues such as relations, proximity and co-movement – semantically complex signals that go beyond traditional object detection. To address this challenge, we introduce a social group region detection task, which requires inferring and spatially grounding visual regions defined by abstract interpersonal relations. We propose MINGLE (Modeling INterpersonal Group-Level Engagement), a modular three-stage pipeline that integrates: (1) off-the-shelf human detection and depth estimation, (2) VLM-based reasoning to classify pairwise social affiliation, and (3) a lightweight spatial aggregation algorithm to localize socially connected groups. To support this task and encourage future research, we present a new dataset of 100K urban street-view images annotated with bounding boxes and labels for both individuals and socially interacting groups. The annotations combine human-created labels and outputs from the MINGLE pipeline, ensuring semantic richness and broad coverage of real world scenarios.
1 Introduction
Object detection is a foundational task in computer vision, and state-of-the-art models, like YOLO and Grounding DINO, excel at localizing discrete, well-defined objects, typically described by simple nouns or phrases [1, 2, 3, 4]. However, an increasing number of real-world applications require the detection of regions defined by semantically complex or abstract criteria, such as interactions of objects, social behaviors, or environmental context. Such tasks require a deeper understanding of context, relationships, and semantics than conventional object detection models are equipped to handle.
@@ 1 Introduction
From an urban design perspective, detecting social groups in public spaces has long been a key focus in pedestrian observational studies. Seminal works by Whyte [5], Gehl [6], and Jacobs [7, 8] emphasize that observing how people gather, linger, and interact in cities is crucial to understanding urban vitality. These studies go far beyond simply counting individuals: Inferring social groups is central to evaluating how built environments support or inhibit human interaction. Such knowledge informs the creation of healthier, more inclusive, and socially vibrant urban environments.
@@
Using MINGLE, we construct a new dataset of 100K images with annotations for both individual people and social groups, making it the first large-scale dataset focused on semantically complex social interaction regions in urban outdoor scenes not limited to a particular interaction type. Our labels combine high-quality manual annotations and carefully validated pseudo-labels propagated through our pipeline.
@@
2 Related Work
2.1 Social Interaction Research
@@
Detecting social groups has long been a key focus in pedestrian observational studies. Starting from classic works by Whyte, Gehl, and Jacobs, more recent studies in urban planning revisit criteria of public space liveliness and emphasize that social interactions play a central role [5, 7, 8, 6, 9, 10].
@@
However, conducting large-scale analyses of group-level social behavior has been limited by the labor- and cost-intensive nature of manual fieldwork. Advances in AI and computer vision are now addressing this challenge, offering tools to automate the detection and analysis of social dynamics at scale. While significant progress has been made in human–human interaction detection, most computer vision approaches focus on relatively constrained scenarios. With existing street-view benchmarks being limited in scale [11], large-scale datasets such as JRDB [12, 13] and AVA [14] primarily support action or gesture recognition in surveillance or indoor contexts. Others target pairwise or small-group social signal understanding, often framed as classification problems based on closed-set of body pose, gaze, or proximity cues [15, 16]. Although these datasets and models offer valuable insights, they do not generalize to open-world, urban street environments, where social formations are more ambiguous and embedded in complex spatial and visual context.
2.2 Open-Vocabulary Object Detection
Figure 2: The result of OVD for social group detection
@@ 2.2 Open-Vocabulary Object Detection
Language-based object detection expands the fixed categories of traditional closed-set detection to free-form descriptions. Current commonly used open-vocabulary object detection and segmentation models, such as YOLO-World, Grounding-DINO, and Grounded SAM, have demonstrated reliable performance on standard object detection or segmentation benchmarks [4, 1, 17]. However, reliance on BERT-based language models limits their ability to comprehend semantically complex language structures. As a result, OVD models struggle to interpret abstract natural language instructions for precise visual grounding, particularly when handling complex instructions and visual scenarios. In particular for social group region detection, they often fail to interpret and reason between complex inter-group relationships and surrounding urban context (Figure 2).
2.3 VLMs for Localization, Object Detection, and Segmentation
@@
Large language models (LLMs) and vision-language models (VLMs) excel at understanding semantically complex commands and visual contexts. However, most general-purpose VLMs lack built-in support for precise spatial tasks such as segmentation and object detection, mostly due to the lack of such training data and specialized training procedures. Recent studies confirm that a transformer architecture trained with a sufficiently large amount of localization data can achieve superior performance on object detection and segmentation tasks [18].
@@
The literature highlights several key limitations of current VLMs in region-level localization tasks. Most generalist VLMs output text-based responses without explicit or reliable spatial grounding, limiting their direct applicability to detection or segmentation. Specialized adaptations have emerged, including learnable coordinate tokens and decoupled localization pipelines, e.g., ROD-MLLM [19]), designed to better align textual instructions with precise visual regions.
@@
Moreover, models such as LISA [20] and u-LLaVA [21] introduce segmentation tokens as prompts within VLMs to improve mask prediction, yet these often still focus on single-object instances, lacking robust mechanisms for multi-object or group-level reasoning. Subsequent approaches such as GLaMM [22], PerceptionGPT [23], and PixelLM [24] extend this with condition tokens for multi-object grounding, but their effectiveness degrades with increasing instruction complexity and relational reasoning demands.
2.4 Semantically Complex Object Detection and Segmentation
@@
Figure 3: Illustration of the three-stage pipeline for detecting semantically complex social interaction regions
@@
A recent review of VLM usage for object detection and segmentation underscores that despite progress achieved through multi-modal pretraining, semantically complex region detection—especially for abstract or relational concepts—remains an open challenge [25]. Their analysis reveals persistent difficulties in grounding abstract queries, rejecting false positives, and accurately localizing regions defined by high-level semantics rather than explicit category labels.
@@
To address this, recent work such as Ground-V [26] has taken a step forward by explicitly training VLMs to align natural language descriptions—including complex and compositional instructions—with pixel-level regions. Ground-V demonstrates improved grounding for instructions involving actions and scene context, but it primarily targets synthetic or simple compositional examples and lacks evaluation in socially grounded, real-world scenarios.
@@
As existing methods fail when regions are described using semantically rich or socially embedded concepts, there is a clear need for methods that extend VLMs toward reasoning-driven localization in complex real-world scenes. Our work fills this gap by focusing on social interaction regions in urban street scenes—an example of semantically complex region detection with real-world societal impact. By combining pairwise reasoning combined with spatial aggregation, and introducing the MINGLE dataset for detailed benchmarking, we contribute toward enabling robust detection of abstract, relational visual concepts that current models struggle to handle as well as to urban planing analysis, enabling the design of healthier, more inclusive cities.
3 Method
@@
This task is characterized by three key challenges 1) the complexity of the text query and visual context; 2) multiple regions that satisfy criteria simultaneously, and 3) the potential for hallucinated references of regions not present in the scene. In the following section, we describe a three-stage workflow that enables overcoming these challenges to efficiently detect socially affiliated groups of people in street-view images (see Figure 3).
3.1 Social Group Region Detection Task
@@
denote, respectively, the set of detected single person bounding boxes and socially affiliated group bounding boxes in image $I_{m}$, where each $g^{m}_{i}$ for image $I_{m}$ contains at least two individual detections from $\mathcal{G}^{m}:|\{p^{m}_{j}\in\mathcal{P}^{m}\mid p^{m}_{j}\subseteq g^{m}_{i}\}|\geq 2$
@@
In order to ensure that pipeline results do not suffer from hallucinated references or struggle with multi-region detections, as common for VLM-based object detection methods [26], we propose a three-stage pipeline leveraging existing object-detection models capabilities.
3.2 Three-Stage Pipeline
The proposed pipeline combines off-the-shelf human detection, VLM-based pairwise reasoning, and a lightweight spatial aggregation algorithm to detect socially affiliated groups in street-view images (see Figure  3).
@@ 3.2 Three-Stage Pipeline
In order to address two common critical issues faced by VLMs applied to object detection—hallucinated references and multiple detections, we design our pipeline based on a robust, high-precision and recall human detector.
@@
We instantiate $\mathcal{D}$ with the ATSS-Swin-L-DyHead model [27]. However, Stage 1 is detector-agnostic: $\mathcal{D}$ may be replaced by any modern human detector, e.g., a YOLOv8 variant [3], for further improvements without requiring any modifications to the downstream pipeline.
@@
For every unordered pair of bounding boxes $(p^{(m)}_{i},p^{(m)}_{j})$, we fine-tune Qwen2-VL-7B and Qwen2.5-VL-7B to perform classification (“Yes”, “No”, or “Not sure”) indicating whether the pair belongs to the same social group using a human-annotated dataset (see Dataset Construction section).
@@
then crop the identical region from (i) the RGB image $I_{m}$ and (ii) the predicted depth map $D_{m}$ obtained with EVP [28]. These two crops,
@@
We initialize each person as a singleton cluster, then iteratively merge the pair of clusters $(C_{a},C_{b})$ that maximally increases an agreement score. The score is defined based on whether merging the clusters aligns with the pairwise annotations in $J^{(m)}$. If $(p^{(m)}_{i},p^{(m)}_{k})=\text{"Yes"}$, then placing $i$ and $k$ in the same cluster increases the score; if the answer is ”No” or ”Not sure” - otherwise.
@@
4 Experiments
4.1 Fine-Tuning Dataset Construction
To construct the fine-tuning dataset, we collected 79,265 human-labeled annotations for pairwise social interactions between single people in the images. In each training image, exactly two individuals are highlighted with bounding boxes. Human annotators were instructed to identify whether the indicated individuals belong to the same social group (“yes”), are independent from each other (“no”) or it is not possible to determine (“not sure”). To ensure that the resulting dataset is balanced with regard to “yes” and “no” answers, the annotation process consisted of two iterations: (1) Human annotators received 46K randomly selected street-view images. (2) Based on the analysis of the first iteration (skewed towards “no” answers), we pre-selected pairs with individuals potentially belonging to the same group based on the first-stage fine-tuned Qwen-2 (7B) model answers. The resulting dataset distribution is described in Table 1.
Split Pairs Yes No Not Sure
@@ 4.1 Fine-Tuning Dataset Construction
Table 1: Pairwise Relationships Dataset Overview
For the model training, the pairwise-relationships dataset is additionally split into different subsets. Variants differ by dataset sampling (balanced vs. unbalanced) and pre-processing (deletion or replacement of low-confidence pairs, i.e., those labeled as “not sure”, where “YN” indicates that all “not sure” answers were removed).
@@
Qwen2-VL 7B LoRA (rank=64) 29k-YN 0.64 0.57 0.56 0.69 0.75 0.49
Qwen2.5-VL 7B LoRA (rank=64) 29k-YN 0.65 0.53 0.55 0.72 0.82 0.42
Qwen2.5-VL 7B LoRA (rank=128) 29k-YN 0.66 0.39 0.52 0.78 0.91 0.26
@@
Qwen2-VL 7B LoRA (rank=64) 61k-Balanced 0.68 0.62 0.59 0.75 0.80 0.53
Qwen2-VL 7B LoRA (rank=64) 76k-Unbalanced 0.70 0.72 0.67 0.75 0.73 0.69
Qwen2-VL 7B LoRA (rank=64) 61k-Balanced-YN 0.72 0.70 0.65 0.78 0.79 0.63
Qwen2-VL 7B LoRA (rank=64) 76k-Unbalanced-YN 0.68 0.65 0.61 0.75 0.78 0.57
@@
Table 2: Model Performance on pairwise relationship classification task. Fine-tuning results in the upper part of the table, zero-shot VLM performance in the lower part
4.2 Model Selection
To identify a strong yet efficient model for pairwise social affiliation classification, we benchmarked several open-source and commercial VLMs on a curated set of 2000 manually validated image pairs. As shown in Table 2, larger models like Qwen2-VL (72B) outperform others in overall $F_{1}$. However, smaller variants—especially Qwen2-VL (7B)—show just marginally lower performance, while being more robust to overfitting, have lower memory requirements and significantly faster inference. The latter is crucial for urban research [^1] Urban analysis requires working with vast amounts of data, e.g. for New-York city alone amount of street-views exceeds 8M., therefore, Qwen2-VL (7B) and Qwen2.5-VL (7B) were selected for further fine-tuning.
4.3 Prompt Generation
To create suitable text prompts for guiding the model towards social group understanding, we evaluated a variety of formulations augmented with additional visual and textual cues (e.g. bounding box coordinates for individual people, numeric depth and distance values, etc.). We determined that the best results are achieved when the model receives inputs and prompt as illustrated in Figure 3. We release nine best-performing prompt versions in the project repository.
4.4 Evaluation
@@
To evaluate the quality of group region detections produced by our pipeline, we adopt the ELSA benchmark [11], as it provides the most relevant content—urban street-view imagery annotated with social group bounding boxes across a range of activities and group sizes.
@@
4.5 Implementation Details
We tested several fine-tuning configurations, including LoRA adapters with rank 64 and 128, and full fine-tuning. For the fine-tuning, we used 8 NVIDIA A100 GPUs and repeated all experiments twice. Each model was trained for 5 epochs and the best-performing one is selected based on the validation set performance.
Figure 4: Descriptive statistics pertaining to our Social Group Region dataset
5 Results
5.1 Fine-tuning

### v3

@@ Abstract
Understanding group-level social interactions in public spaces is crucial for urban planning, informing the design of socially vibrant and inclusive environments. Detecting such interactions from images involves interpreting subtle visual cues such as relations, proximity, and co-movement—semantically complex signals that go beyond traditional object detection. To address this challenge, we introduce a social group region detection task, which requires inferring and spatially grounding visual regions defined by abstract interpersonal relations. We propose MINGLE (Modeling INterpersonal Group-Level Engagement), a modular three-stage pipeline that integrates: (1) off-the-shelf human detection and depth estimation, (2) VLM-based reasoning to classify pairwise social affiliation, and (3) a lightweight spatial aggregation algorithm to localize socially connected groups. To support this task and encourage future research, we present a new dataset of 100K urban street-view images annotated with bounding boxes and labels for both individuals and socially interacting groups. The annotations combine human-created labels and outputs from the MINGLE pipeline, ensuring semantic richness and broad coverage of real-world scenarios.
Code and Datasethttps://github.com/lyons66/MINGLE
Introduction
Object detection is a foundational task in computer vision, and state-of-the-art models, such as YOLO and Grounding DINO, excel at localizing discrete, well-defined objects, typically described by simple nouns or phrases (Liu et al. 2024; Ren et al. 2016; Jocher et al. 2023; Cheng et al. 2024). However, an increasing number of real-world applications require the detection of regions defined by semantically complex or abstract criteria, e.g., interactions of objects, social behaviors, or environmental context. Such tasks require a deeper understanding of context, relationships, and semantics than conventional object detection models are equipped to handle.
@@ Introduction
From an urban design perspective, detecting social groups in public spaces has long been a key focus in pedestrian observational studies. Seminal works by Whyte (Whyte and others 1980), Gehl (Gehl 1987), and Jacobs (Jacobs 1961, 1985) emphasize that observing how people gather, linger, and interact in cities is crucial to understanding urban vitality. These studies go far beyond simply counting individuals: Inferring social groups is central to evaluating how built environments support or inhibit human interaction. Such knowledge informs the creation of healthier, more inclusive, and socially vibrant urban environments.
@@
Using MINGLE, we construct a new dataset of 100K images with annotations for both individual people and social groups, making it the first large-scale dataset focused on semantically complex social interaction regions in urban outdoor scenes, not limited to a particular interaction type. Our labels combine high-quality manual annotations and carefully validated pseudo-labels propagated through our pipeline.
@@
Related Work
Social Interaction Research
@@
Detecting social groups has long been a key focus in pedestrian observational studies. Starting from classic works by Whyte, Gehl, and Jacobs, more recent studies in urban planning revisit criteria of public space liveliness and emphasize that social interactions play a central role (Whyte and others 1980; Jacobs 1961, 1985; Gehl 1987; Mehta and Bosson 2021; Qi et al. 2024).
@@
However, conducting large-scale analyses of group-level social behavior has been limited by the labor- and cost-intensive nature of manual fieldwork. Advances in AI and computer vision aim to address this challenge, offering tools to automate the detection and analysis of social dynamics at scale. Despite significant progress in human–human interaction detection, most computer vision approaches focus on relatively constrained scenarios. With existing street-view benchmarks being limited in scale (Hosseini et al. 2024), large-scale datasets such as JRDB (Ehsanpour et al. 2022; Jahangard et al. 2024) and AVA (Gu et al. 2018) primarily support action or gesture recognition in surveillance or indoor contexts. Others target pairwise or small-group social signal understanding, often framed as classification problems based on a closed set of body pose, gaze, or proximity cues (Wang et al. 2023; Wei et al. 2024). Although these datasets and models offer valuable insights, they do not generalize to open-world, urban street environments, where social formations are more ambiguous and embedded in complex spatial and visual contexts.
Open-Vocabulary Object Detection
Figure 2: The result of OVD for social group detection.
@@
Language-based object detection expands the fixed categories of traditional closed-set detection to free-form descriptions. Current commonly used open-vocabulary object detection and segmentation models, such as YOLO-World, Grounding-DINO, and Grounded SAM, have demonstrated reliable performance on standard object detection or segmentation benchmarks (Cheng et al. 2024; Liu et al. 2024; Ren et al. 2024a). However, reliance on BERT-based language models limits their ability to comprehend semantically complex language structures. As a result, OVD models struggle to interpret abstract natural language instructions for precise visual grounding, particularly when handling complex instructions and visual scenarios. In particular, for social group region detection, they often fail to interpret and reason between complex inter-group relationships and the surrounding urban context (Figure 2).
VLMs for Localization, Object Detection, and Segmentation
@@
Large language models (LLMs) and vision-language models (VLMs) excel at understanding semantically complex commands and visual contexts. However, most general-purpose VLMs lack built-in support for precise spatial tasks such as segmentation and object detection, mostly due to the lack of such training data and specialized training procedures. Recent studies confirm that a Transformer architecture trained with a sufficiently large amount of localization data can achieve superior performance on object detection and segmentation tasks (Xiao et al. 2023).
@@
The literature highlights several key limitations of current VLMs in region-level localization tasks. Most generalist VLMs output text-based responses without explicit or reliable spatial grounding, limiting their direct applicability to detection or segmentation. Specialized adaptations have emerged, including learnable coordinate tokens and decoupled localization pipelines, e.g., ROD-MLLM (Yin et al. 2025), designed to better align textual instructions with precise visual regions.
@@
Moreover, models such as LISA (Lai et al. 2024) and u-LLaVA (Xu et al. 2024) introduce segmentation tokens as prompts within VLMs to improve mask prediction, yet these often still focus on single-object instances, lacking robust mechanisms for multi-object or group-level reasoning. Subsequent approaches such as GLaMM (Rasheed et al. 2024), PerceptionGPT (Pi et al. 2023), and PixelLM (Ren et al. 2024b) extend this with condition tokens for multi-object grounding, but their effectiveness degrades with increasing instruction complexity and relational reasoning demands.
Semantically Complex Object Detection and Segmentation
@@
Figure 3: Illustration of the three-stage pipeline for detecting semantically complex social interaction regions.
@@
A recent review of VLM usage for object detection and segmentation underscores that despite progress achieved through multi-modal pretraining, semantically complex region detection—especially for abstract or relational concepts—remains an open challenge (Feng et al. 2025). Their analysis reveals persistent difficulties in grounding abstract queries, rejecting false positives, and accurately localizing regions defined by high-level semantics rather than explicit category labels.
@@
To address this, recent work such as Ground-V (Zong et al. 2025) has taken a step forward by explicitly training VLMs to align natural language descriptions—including complex and compositional instructions—with pixel-level regions. Ground-V demonstrates improved grounding for instructions involving actions and scene context, but it primarily targets synthetic or simple compositional examples and lacks evaluation in socially grounded, real-world scenarios.
@@
As existing methods fail when regions are described using semantically rich concepts, there is a clear need for methods that extend VLMs toward reasoning-driven localization in complex real-world scenes. Our work fills the gap by focusing on social interaction regions in urban scenes—an example of semantically complex region detection with real-world societal impact. By combining pairwise reasoning with spatial aggregation and introducing the MINGLE dataset, we contribute toward enabling robust detection of abstract, relational visual concepts that current models struggle to handle, as well as to urban planning analysis, enabling the design of healthier, more inclusive cities.
Method
@@
This task is characterized by three key challenges: 1) the complexity of the text query and visual context; 2) multiple regions that satisfy criteria simultaneously, and 3) the potential for hallucinated references of regions not present in the scene. In the following section, we describe a three-stage workflow that enables overcoming these challenges to efficiently detect socially affiliated groups of people in street-view images (see Figure 3).
Social Group Region Detection Task
@@
denote, respectively, the set of detected single person bounding boxes and socially affiliated group bounding boxes in image $I_{m}$, where each $g^{m}_{i}$ for image $I_{m}$ contains at least two individual detections from $\mathcal{G}^{m}$ such that $\{p^{m}_{j}\in\mathcal{P}^{m}\mid p^{m}_{j}\subseteq g^{m}_{i}\}|\geq 2$.
@@
In order to ensure that pipeline results do not suffer from hallucinated references or struggle with multi-region detections, as is common for VLM-based object detection methods (Zong et al. 2025), we propose a three-stage pipeline leveraging existing object-detection models’ capabilities.
Three-Stage Pipeline
The proposed pipeline combines off-the-shelf human detection, VLM-based pairwise reasoning, and a lightweight spatial aggregation algorithm to detect socially affiliated groups in street-view images (see Figure 3).
@@
In order to address two common critical issues faced by VLMs applied to object detection—hallucinated references and multiple detections, we design our pipeline based on a robust, high-precision, and recall human detector.
@@
We instantiate $\mathcal{D}$ with the ATSS-Swin-L-DyHead model (Zhang et al. 2020). However, Stage 1 is detector-agnostic: $\mathcal{D}$ may be replaced by any modern human detector, e.g., a YOLOv8 variant (Jocher et al. 2023), for further improvements without requiring any modifications to the downstream pipeline.
@@
For every unordered pair of bounding boxes $(p^{(m)}_{i},p^{(m)}_{j})$, we fine-tune Qwen2-VL-7B and Qwen2.5-VL-7B to perform classification (“Yes”, “No”, or “Not sure”), indicating whether the pair belongs to the same social group using a human-annotated dataset (see Dataset Construction section).
@@
then crop the identical region from (i) the RGB image $I_{m}$ and (ii) the predicted depth map $D_{m}$ obtained with EVP (Lavreniuk et al. 2024). These two crops,
@@
We initialize each person as a singleton cluster, then iteratively merge the pair of clusters $(C_{a},C_{b})$ that maximally increases an agreement score. The score is defined based on whether merging the clusters aligns with the pairwise annotations in $J^{(m)}$. If $(p^{(m)}_{i},p^{(m)}_{k})=\text{"Yes"}$, then placing $i$ and $k$ in the same cluster increases the score, while if the answer is ”No” or ”Not sure”, it does not.
@@
Experiments
Fine-Tuning Dataset Construction
To construct the fine-tuning dataset, we collected 79,265 human-labeled annotations for pairwise social interactions between single people in the images. In each training image, exactly two individuals are highlighted with bounding boxes. Human annotators were instructed to identify whether the indicated individuals belong to the same social group (“yes”), are independent from each other (“no”), or it is not possible to determine (“not sure”). To ensure that the resulting dataset is balanced with regard to “yes” and “no” answers, the annotation process consisted of two iterations: (1) Human annotators received 46K randomly selected street-view images. (2) Based on the analysis of the first iteration (skewed towards “no” answers), we pre-selected pairs with individuals potentially belonging to the same group based on the first-stage fine-tuned Qwen-2 (7B) model answers. The resulting dataset distribution is described in Table 1.
Split Pairs ”Yes” ”No” ”Not Sure”
@@
Table 1: Pairwise Relationships Dataset Overview.
For the model training, the pairwise-relationships dataset is additionally split into different subsets. Variants differ by dataset sampling (balanced vs. unbalanced) and pre-processing (deletion or replacement of low-confidence pairs, i.e., those labeled as “not sure”, where “YN” indicates that all “not sure” answers were omitted).
@@
Qwen2-VL 7B LoRA (rank 64) 29k-YN 0.64 0.57 0.56 0.69 0.75 0.49
Qwen2.5-VL 7B LoRA (rank 64) 29k-YN 0.65 0.53 0.55 0.72 0.82 0.42
Qwen2.5-VL 7B LoRA (rank 128) 29k-YN 0.66 0.39 0.52 0.78 0.91 0.26
@@
Qwen2-VL 7B LoRA (rank 64) 61k-Balanced 0.68 0.62 0.59 0.75 0.80 0.53
Qwen2-VL 7B LoRA (rank 64) 76k-Unbalanced 0.70 0.72 0.67 0.75 0.73 0.69
Qwen2-VL 7B LoRA (rank 64) 61k-Balanced-YN 0.72 0.70 0.65 0.78 0.79 0.63
Qwen2-VL 7B LoRA (rank 64) 76k-Unbalanced-YN 0.68 0.65 0.61 0.75 0.78 0.57
@@
Table 2: Model Performance on pairwise relationship classification task. Fine-tuning results in the upper part of the table, zero-shot VLM performance in the lower part.
Model Selection
To identify a strong yet efficient model for pairwise social affiliation classification, we benchmarked several open-source and commercial VLMs on a curated set of 2,000 manually validated image pairs. As shown in Table 2, larger models like Qwen2-VL (72B) outperform others in overall $F_{1}$. However, smaller variants—especially Qwen2-VL (7B)—obtain just marginally lower effectiveness, while offering greater robustness to overfitting, lower memory consumption, and significantly faster inference. The latter is crucial for urban research, [^1] Urban analysis requires working with vast amounts of data; e.g., for New York City alone, the number of street views exceeds 8M. and therefore Qwen2-VL (7B) and Qwen2.5-VL (7B) were selected for further fine-tuning.
Prompt Generation
To create suitable text prompts for guiding the model towards social group understanding, we evaluated a variety of formulations augmented with additional visual and textual cues (e.g., bounding box coordinates for individual people, numeric depth and distance values, etc.). We determined that the best results are achieved when the model receives inputs and prompts as illustrated in Figure 3. We release the nine best-performing prompt versions in the project repository.
Evaluation
@@
To evaluate the quality of group region detections produced by our pipeline, we adopt the ELSA benchmark (Hosseini et al. 2024), as it provides the most relevant content—urban street-view imagery annotated with social group bounding boxes across a range of activities and group sizes.
@@
Implementation Details
We tested several fine-tuning configurations, including LoRA adapters with rank 64 and 128, and full fine-tuning. For fine-tuning, we used 8 NVIDIA A100 GPUs and repeated all experiments twice. Each model was trained for 5 epochs and the best-performing one is selected based on the validation set performance.
Figure 4: Descriptive statistics pertaining to our Social Group Region dataset.
Results
Fine-tuning